OpenStudy (mathhelp346):

The product of a certain number plus 3 and that same number minus 5 is 9. Set up an equation and solve the equation to find the number.

OpenStudy (mathhelp346):

my equation was (x+3)(x-5)=9 but i don't know if that's right

OpenStudy (mathhelp346):

i got x=16

OpenStudy (oaktree):

(x+3)(x-5) = 9 x^2-2x-15=9 x^2-2x-24=0 (x+4)(x-6)=0 x=-4,6

OpenStudy (mathhelp346):

ohhhh ok

OpenStudy (mathhelp346):

thanks

OpenStudy (mathhelp346):

@OakTree why did u distribute at the first step?

OpenStudy (oaktree):

That's because you can't solve for the roots until your equation is equal to zero. Basic rule of algebra!